The probablility is 1 in 4,000,000 that a single auto trip in the United States will result in a fatality. Over a lifetime, an average US driver takes 50,000 trips. (a) What is the probablilty of a fatal accident over a lifetime? Explain your reasoning carefully. Hint: Assume independent events. Why might the assumption of independence be violated? (b) Why might a driver be tempted not to use a seat belt "just on this trip?"
Answer by checkley71(8403)

1/4,000,000=X/50,000
4,000,000X=50,000
X=50,000/4,000,000
X=.0125 OR 1.25% CHANCE OF A FATALITY IN 50,000 MILES DRIVEN.